OpenStudy (anonymous):

Provide the complete solution for the following linear congruences. 3x-2 ≡ 7 (mod 11) 5 ≡ 4x-1 (mod 9) Provide the complete solution for the following linear congruences. 3x-2 ≡ 7 (mod 11) 5 ≡ 4x-1 (mod 9) @Mathematics

OpenStudy (anonymous):

what does "a complete solution" look like?

OpenStudy (anonymous):

first one is \[x\equiv 3(\text{ mod }11)\] and second is \[x\equiv 6 ( \text { mod } 9)\] so you need to solve them both?

OpenStudy (anonymous):

well for x ≡ 3 (mod 7) x ≡ 5 (mod 12) the solution is x ≡ 17 (mod 84)

OpenStudy (anonymous):

it's a system of linear congruences so the x has to be for both

OpenStudy (anonymous):

??

OpenStudy (anonymous):

where does the 17 come from?

OpenStudy (anonymous):

here is an example solution: n ≡ 5 (mod 12) n ≡ 11 (mod 18) The first congruence implies n = 5+12x for some x 2 Z. Substitute this into the second congruence to get 5 + 12x ≡ 11 (mod 18), or 12x ≡ 6 (mod 18). Since gcd(12, 18) = 6 which divides 6, there is a solution. We can use EEA or by inspection to see that x0 = 2 is one solution, hence by LCT 1, the complete solution to the second congruence is x ≡ 2 (mod 18/6) or x ≡ 2 (mod 3). So x = 2 + 3y for some y 2 Z. Substitute back into n to get n = 5 + 12(2 + 3y) = 29 + 36y. So the complete solution is n ≡ 29 (mod 36).

OpenStudy (anonymous):

those x 2 Z should be x is and element of the integers and y is an element of the integers

OpenStudy (anonymous):

ok i think the first equation gives \[x\equiv 3(\text{ mod }11)\] and the second gives \[x\equiv 6 ( \text { mod } 9)\] does this look reasonable? then we can solve as above

OpenStudy (anonymous):

I will show you what I've got so far, give me a second

OpenStudy (anonymous):

\[x=3+11k\] \[3+11k\equiv 6(\text { mod } 9)\] \[11k\equiv 3 (\text { mod } 9)\] and by inspection \[k=6\] works since \[66\equiv 3( \text{ mod } 9)\]

OpenStudy (anonymous):

http://dl.dropbox.com/u/6717478/part.png is what I have so far

OpenStudy (anonymous):

it looks like you might have made a mistake in the line that says \[5\equiv 4x-1 (\text { mod }11)\] \[4\equiv 4x( \text { mod }11)\] shouldn't this be \[6\equiv 4x( \text { mod }11)\]

OpenStudy (anonymous):

oh you're right

OpenStudy (anonymous):

then \[4x\equiv 6 ( mod 9)\] \[2x\equiv 3( mod 9)\] \[2x\equiv 12(mod 9)\] \[x\equiv 6 (mod 9)\] or did i make a mistake there? i think this is right,

OpenStudy (anonymous):

but you should be able to proceed with your method once you correct that error right?

OpenStudy (anonymous):

I will try, when I was writing it out I started working more of it out

OpenStudy (anonymous):

good luck!

OpenStudy (anonymous):

thanks for your help!

OpenStudy (anonymous):

I think the answer is complete solution is x ≡ 69 (mod 99)

OpenStudy (anonymous):

i have what i think may be right but you can correct me if i am wrong. i want to solve \[x\equiv 3(11)\] and \[x\equiv 6(9)\] so i write \[x=3+11k\] and then \[3+11k\equiv 6(9)\] \[11k\equiv 3(9)\] and then i keep adding 9 until i can divide by 11 and get \[11k\equiv 66(9)\] and so \[k\equiv 3(9)\] so \[k=3+9j\] and substitute back to get \[x = 11(3+9j)+3\] \[x=36+99j\] and therefore \[x\equiv 36(99)\] but don't take my word for this

OpenStudy (anonymous):

too bad i got 66 and you got 69, because now i doubt my method. let me check my arithmetic

OpenStudy (anonymous):

no mine is wrong

OpenStudy (anonymous):

in the second equation if you put int 36 then you get 5 ≡ 4(36)-1 (mod 9) 5 ≡ 143 (mod 9) 5 ≡ 8 (mod 9) which isn't true I don't think

OpenStudy (anonymous):

no you are right and i am wrong. i am not sure why but i will try to figure it out

OpenStudy (anonymous):

whatever you did worked when i checked for 69 + 99

OpenStudy (anonymous):

ok that's a relief haha I've been stuck on this one for a while

OpenStudy (anonymous):

i don't want you to do more work, but isn't \[5\equiv 4x-1(9)\] the same as \[x\equiv 6(9)\]?

OpenStudy (anonymous):

never mind you are done and i will figure it out on my own

OpenStudy (anonymous):

what did you do with the 4?

OpenStudy (anonymous):

i put \[6\equiv 4x(9)\] \[4x\equiv 6(9)\] then either divide by 2 and solve or add 9 until i can divide by 4 i get \[4x\equiv 24(9)\] and so \[x\equiv 6(9)\]

OpenStudy (anonymous):

5 ≡ 4x-1 (mod 9) 6 ≡ 4x (mod 9) 6 - 4x ≡ 0 (mod 9) -4x ≡ -6 (mod 9) I don't know if you can multiply the - away

OpenStudy (anonymous):

you can

OpenStudy (anonymous):

oh I don't think you can do division in the linear congruence or else something like 13x ≡ 1 (mod 60) would be very trivial wouldn't it?

OpenStudy (anonymous):

or simply switch them \[6\equiv 4x(9)\] \[4x\equiv 6(9)\]

OpenStudy (anonymous):

I don't think you can divide by the 4 to get rid of it like that

OpenStudy (anonymous):

equivalence is an equivalence relation, so i can switch.

OpenStudy (anonymous):

at least I've never seen my prof do that so I never have, but I'm not sure

OpenStudy (anonymous):

ok well i will figure out my mistake, i will not take any more of your time. good work though!

OpenStudy (anonymous):

thanks, and thank you for always helping me! I may need some more help later on if I get stuck haha